OpenStudy (anonymous):

the equation of the line that passes through the point (–2, 1) and has slope 3 in point-slope form.

OpenStudy (katrinakaif):

Plug into the equation: (y-y1) = m (x-x1) (y-1) = 3(x-(-2) and Solve